convert 11011 base2 to a number in base 10
Answer by stanbon(75887)

convert 11011 base2 to a number in base 10
------
From left to right you have: 2^4+2^3+2+1 = 16+8+3 = 27
====================================================
Cheers,
Stan H.
==============